The equation Qd=1000-20p;work out the quantity demanded for price of RS 5,10,20 and 50
1
Expert's answer
2020-12-23T07:18:02-0500

Qd for price 5 is equal to 900 (it can be calculated as 1000 - 20 "\ltimes" 5 = 1000 - 100 = 900).

Qd for price 10 is equal to 900 (it can be calculated as 1000 - 20 "\ltimes" 10 = 1000 - 200 = 800).

Qd for price 20 is equal to 900 (it can be calculated as 1000 - 20 "\ltimes" 20 = 1000 - 400 = 600).

Hence,

Qd = 900 when p=5

Qd = 800 when p=10

Qd = 600 when p=20

In accordance with the law of demand and equation Qd=1000-20 "\ltimes" p, conditional on all else being equal, as the price of a good goes up (from 5 to 20), quantity demanded will go down (from 900 to 600).